OpenStudy (anonymous):

please explain how to do this !! "For each given pair of functions, multiply or divide as instructed. Be sure to show all the steps required to write your answer in simplest form. " f(x) = 2x g(x) = 3x^3 i know how to combine the two, i think ? im guessing that combined they are 2x * (3x^3). correct me if im wrong, and please help me distribute and simplify !

Nnesha (nnesha):

when you multiply same variables you should add their exponents \[\huge\rm x^m \times x^n = x^{ m \times n }\]

Nnesha (nnesha):

f(x) times g(x) replace f(x) and g(x) by their values then multiply

OpenStudy (anonymous):

okay ... so i multiply 2x and 3x^3 ? that would be 6^3, right ?

Nnesha (nnesha):

6^3 ?? 6 power 3 ?

Nnesha (nnesha):

when you multiply variables with SAME bases then add their exponent like in this exponent rule x^m times x^n both bases are same so x^{m+n}

Nnesha (nnesha):

you can ask question if you don't get it :-)

OpenStudy (anonymous):

sorry i had to do something else for my mom and then i was trying to figure it out. i dont quite understand yet though. what do i do with the coefficient ?

Nnesha (nnesha):

multiply them together 2x times 3x^3 2 times 3 x times x^3

OpenStudy (anonymous):

oooh okay !!! so, 2 * 3 = 6. and x * x^3 = x^3, right ? now do i combine those answers ?

Nnesha (nnesha):

like i said when you multiply same variables you should add their exponent and remember x is raising to the one power

Nnesha (nnesha):

\[\large\rm x^1 \times x^3 \] apply exponent rule

OpenStudy (anonymous):

oh right, sorry :PP i completely forgot that, haha so i wouldnt 'multiply' per say, id add, and it would be x^4 ??

Nnesha (nnesha):

yep!

OpenStudy (anonymous):

great, okay ! is the next step to combine the answers ?

Nnesha (nnesha):

are there any like terms ??

OpenStudy (anonymous):

no, there isnt i dont think ! is that the answer ??

Nnesha (nnesha):

yep

Nnesha (nnesha):

btw after multiplying is there any term left to combine 2 times 3 x^1 times x^3 = x^4 6x^4 i separated these so you can understand :-)

OpenStudy (anonymous):

thank you so much for your help

Nnesha (nnesha):

no problem do you have to divide that ?

OpenStudy (anonymous):

no i dont, thank you though : ))

Nnesha (nnesha):

nice! my pleasure :-)